# Kernel vs. Kernel: Exploring How the Data Structure Affects Neural Collapse

### v1

@@ 5.2 Limiting NTK
$\displaystyle\begin{split}K_{GP}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})&=\sigma_{b}^{2}+\frac{\sigma_{w}^{2}}{d_{1}}Q_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).\end{split}$ (14)
@@
Remark on the limiting NTK. Denote by $w,\phi$ the parameters and the activation function of an $L$-layer NN $\psi(\cdot)$. The limiting NNGP is defined directly on the product of neurons, $\mathbb{E}_{w}\langle\phi(z_{i}(\mathbf{x})),\phi(z_{i}(\tilde{\mathbf{x}}))\rangle$, and provides kernel expressions for the inner product of features at different layers of the network (via recursion similar to (11)). On the other hand, the limiting NTK is defined as the inner product of the output’s gradients $\mathbb{E}_{w}\langle\nabla_{w}\psi(\mathbf{x}),\nabla_{w}\psi(\tilde{\mathbf{x}})\rangle$. The NTK theory shows that this can model the inner product only of the deepest features (i.e the output of the penultimate layer).
@@ Appendix B Limiting NNGP and NTK for ReLU
$\displaystyle\begin{split}K_{GP-ReLU}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})&=\sigma_{b}^{2}+\frac{\sigma_{w}^{2}}{d_{1}}Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).\end{split}$ (26)
@@ D.2 NC1 of limiting NTK with ReLU activation
$\displaystyle\begin{split}K_{GP-ReLU}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})&=\sigma_{b}^{2}+\frac{\sigma_{w}^{2}}{d_{1}}Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).\end{split}$ (50)
@@
$\displaystyle\theta_{c,i}^{c^{\prime},j}=\begin{cases}0&\text{if }c=c^{\prime}\\ \pi&\text{if }c\neq c^{\prime}\\ \end{cases}$ (52)
When $\sigma_{b}\to 0,d_{1}\to\infty$, we get:
$\displaystyle\Theta^{(2)}_{NTK-ReLU}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=K_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})\dot{Q}^{(1)}_{GP-ReLU}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ (53)
$\displaystyle\implies\Theta_{NTK-ReLU}^{(2)}(x^{c,i},x^{c^{\prime},j})$ =$\displaystyle=\begin{cases}\frac{\sigma_{w}^{2}}{2}x^{c,i}x^{c,j}&\text{if }c=c^{\prime}\\ 0&\text{if }c\neq c^{\prime}\end{cases},$ (54)
which resembles (45) for $Q_{GP-ReLU}^{(1)}(x^{c,i},x^{c^{\prime},j})$. Thus, we end up with the same result as (48):
@@ E.2 NC1 of Limiting NTK with Erf activation
$\displaystyle K_{GP-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})$ $\displaystyle=\sigma_{b}^{2}+\frac{\sigma_{w}^{2}}{d_{1}}Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})$ (100)
@@
Considering $\sigma_{b}\to 0$, $d_{0}=1$ and $d_{1}\to\infty$ (as per the setting and assumptions), we get:
@@
$\displaystyle\dot{Q}_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})$ $\displaystyle=\frac{4}{\pi}\left(\left(1+2\sigma_{w}^{2}x^{c,i}x^{c,i}\right)\left(1+2\sigma_{w}^{2}x^{c^{\prime},j}x^{c^{\prime},j}\right)-\left(2\sigma_{w}^{2}x^{c,i}x^{c^{\prime},j}\right)^{2}\right)^{-1/2}$ (104)
$\displaystyle=\frac{4}{\pi\sqrt{1+2\sigma_{w}^{2}\cdot(x^{c,i})^{2}+2\sigma_{w}^{2}\cdot(x^{c^{\prime},j})^{2}}}.$ (105)
@@
$\displaystyle\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})$ $\displaystyle=K_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})\dot{Q}_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})$ (106)
$\displaystyle=\frac{4\sigma_{w}^{2}x^{c,i}x^{c^{\prime},j}}{\pi\sqrt{1+2\sigma_{w}^{2}\cdot(x^{c,i})^{2}+2\sigma_{w}^{2}\cdot(x^{c^{\prime},j})^{2}}}$ (107)
@@
$\displaystyle\kappa(x^{c,i},x^{c^{\prime},j})=\sqrt{\frac{1}{\sigma_{w}^{2}(x^{c,i})^{2}(x^{c^{\prime},j})^{2}}+\frac{2}{(x^{c^{\prime},j})^{2}}+\frac{2}{(x^{c,i})^{2}}},$ (110)
@@
$\displaystyle\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})$ $\displaystyle=\frac{4\sigma_{w}\operatorname{sign}(x^{c,i})\operatorname{sign}(x^{c^{\prime},j})}{\pi\kappa(x^{c,i},x^{c^{\prime},j})}$ (111)
@@ E.2.1 Calculating $\mathbb{E}\left[\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})\right]$
$\displaystyle\begin{split}\mathbb{E}\left[\kappa(x^{c,i},x^{c,i})\right]&=\frac{1}{2}+2T(c)+\frac{1}{2\sigma_{w}^{2}}\left[\frac{1}{3\sigma_{c}^{4}+6\sigma_{c}^{2}\mu_{c}^{2}+\mu_{c}^{4}}\right]+\mathbb{E}[\xi_{h.o.t}]\end{split}$ (117)
@@
$\displaystyle\begin{split}S(\mu_{c},\sigma_{c})&=-\frac{1}{2}+2T(c)+\frac{1}{2\sigma_{w}^{2}}\left[\frac{1}{3\sigma_{c}^{4}+6\sigma_{c}^{2}\mu_{c}^{2}+\mu_{c}^{4}}\right]+\mathbb{E}[\xi_{h.o.t}]\end{split}$ (118)
@@
$\displaystyle\mathbb{E}\left[\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})\right]$ =$\displaystyle=\begin{cases}\mathbb{E}\left[\frac{4\sigma_{w}}{\pi\kappa(x^{c,i},x^{c,j})}\right]&c=c^{\prime}\\ -\mathbb{E}\left[\frac{4\sigma_{w}}{\pi\kappa(x^{c,i},x^{c^{\prime},j})}\right]&c\neq c^{\prime}\end{cases},$ (141)
which can now be expanded into:
@@
=$\displaystyle\hskip 40.0pt=\begin{cases}\frac{4\sigma_{w}}{\pi}\left(1-S(\mu_{c},\sigma_{c})+\widetilde{\delta}_{h.o.t}(\kappa(x^{c,i},x^{c,i}))\right)&c=c^{\prime};i=j\\ \frac{4\sigma_{w}}{\pi}\left(\frac{3}{2}-\frac{T(c)^{2}}{2\sigma_{w}^{2}}-2T(c)+\widetilde{\delta}_{h.o.t}(\kappa(x^{c,i},x^{c,j}))\right),&c=c^{\prime},i\neq j\\ -\frac{4\sigma_{w}}{\pi}\left(\frac{3}{2}-\frac{T(c)T(c^{\prime})}{2\sigma_{w}^{2}}-T(c)-T(c^{\prime})+\widetilde{\delta}_{h.o.t}(\kappa(x^{c,i},x^{c^{\prime},j}))\right),&c\neq c^{\prime}\\ \end{cases}$ (143)
Observe that these cases resemble the NNGP scenario (up to scaling, where the additional scaling factor gets canceled when calculating $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$) in (85) and thus follow similar conclusions regarding the role of $\mu_{c},\sigma_{c}$ in determining the value of $\mathbb{E}\left[\mathcal{N}\mathcal{C}_{1}(\mathbf{H})\right]$.

### v2

@@ 5.2 Limiting NTK
$\displaystyle\begin{split}K_{GP}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=\sigma_{b}^{2}+\sigma_{w}^{2}Q_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).\end{split}$ (14)
@@
Remark on the limiting NTK. Denote by $w,\phi$ the parameters and the activation function of an $L$-layer NN $\psi(\cdot)$. The limiting NNGP is defined directly on the product of neurons, $\mathbb{E}_{w}\langle\phi(z_{i}(\mathbf{x})),\phi(z_{i}(\tilde{\mathbf{x}}))\rangle$, and provides kernel expressions for the inner product of features at different layers of the network (via recursion similar to (11)). On the other hand, the limiting NTK is defined as the inner product of the output’s gradients $\mathbb{E}_{w}\langle\nabla_{w}\psi(\mathbf{x}),\nabla_{w}\psi(\tilde{\mathbf{x}})\rangle$. The NTK theory shows that this can model the inner product only of the deepest features (i.e the output of the penultimate layer) [^1] Note that we denote NNGP $(Q_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}))$ and NTK ($\Theta^{(2)}_{NTK}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$) with different superscripts to be consistent with the literature. Yet both are associated with the output of the single hidden layer. .
@@ Appendix B Limiting NNGP and NTK for ReLU
$\displaystyle\begin{split}K_{GP-ReLU}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=\sigma_{b}^{2}+\sigma_{w}^{2}Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).\end{split}$ (26)
@@ D.2 NC1 of limiting NTK with ReLU activation
$\displaystyle\begin{split}K_{GP-ReLU}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=\sigma_{b}^{2}+\sigma_{w}^{2}Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).\end{split}$ (50)
@@
$\displaystyle\theta_{c,i}^{c^{\prime},j}=\begin{cases}0&\text{if }c=c^{\prime}\\ \pi&\text{if }c\neq c^{\prime}\\ \end{cases}.$ (52)
When $\sigma_{b}\to 0$, we get:
$\displaystyle\Theta^{(2)}_{NTK-ReLU}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=\sigma_{w}^{2}Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})+K_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})\dot{Q}^{(1)}_{GP-ReLU}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}).$ (53)
From (45), we know that:
$\displaystyle\implies\Theta_{NTK-ReLU}^{(2)}(x^{c,i},x^{c^{\prime},j})$ =$\displaystyle=\begin{cases}\frac{\sigma_{w}^{4}}{2}x^{c,i}x^{c,j}+\frac{\sigma_{w}^{2}}{2}x^{c,i}x^{c,j}&\text{if }c=c^{\prime}\\ 0&\text{if }c\neq c^{\prime}\end{cases},$ (55)
=$\displaystyle=\begin{cases}\left(\frac{\sigma_{w}^{4}}{2}+\frac{\sigma_{w}^{2}}{2}\right)x^{c,i}x^{c,j}&\text{if }c=c^{\prime}\\ 0&\text{if }c\neq c^{\prime}\end{cases}.$ (56)
Notice that $\Theta_{NTK-ReLU}^{(2)}(x^{c,i},x^{c^{\prime},j})$ is a scaled version of $Q_{GP-ReLU}^{(1)}(x^{c,i},x^{c^{\prime},j})$ (as per (45)). Thus, we end up with the same result as (48) :
@@ E.2 NC1 of Limiting NTK with Erf activation
$\displaystyle K_{GP-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})$ $\displaystyle=\sigma_{b}^{2}+\sigma_{w}^{2}Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})$ (102)
@@
Considering $\sigma_{b}\to 0$, $d_{0}=1$ (as per the setting and assumptions), we get:
@@
$\displaystyle Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})=\frac{2}{\pi}\arcsin\left(\frac{2\sigma_{w}^{2}x^{c,i}x^{c^{\prime},j}}{\sqrt{1+2\sigma_{w}^{2}(x^{c,i})^{2}}\sqrt{1+2\sigma_{w}^{2}(x^{c^{\prime},j})^{2}}}\right).$ (106)
$\displaystyle\begin{split}\dot{Q}_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})&=\frac{4}{\pi}\left(\left(1+2\sigma_{w}^{2}x^{c,i}x^{c,i}\right)\left(1+2\sigma_{w}^{2}x^{c^{\prime},j}x^{c^{\prime},j}\right)-\left(2\sigma_{w}^{2}x^{c,i}x^{c^{\prime},j}\right)^{2}\right)^{-1/2}\\ &=\frac{4}{\pi\sqrt{1+2\sigma_{w}^{2}\cdot(x^{c,i})^{2}+2\sigma_{w}^{2}\cdot(x^{c^{\prime},j})^{2}}}.\end{split}$ (107)
@@
$\displaystyle K_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})\dot{Q}_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})$ $\displaystyle=\frac{4\sigma_{w}^{2}x^{c,i}x^{c^{\prime},j}}{\pi\sqrt{1+2\sigma_{w}^{2}\cdot(x^{c,i})^{2}+2\sigma_{w}^{2}\cdot(x^{c^{\prime},j})^{2}}}$ (108)
@@
$\displaystyle\kappa(x^{c,i},x^{c^{\prime},j})=\sqrt{\frac{1}{\sigma_{w}^{2}(x^{c,i})^{2}(x^{c^{\prime},j})^{2}}+\frac{2}{(x^{c^{\prime},j})^{2}}+\frac{2}{(x^{c,i})^{2}}}$ (111)
@@
$\displaystyle\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})=\sigma_{w}^{2}Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})+\frac{4\sigma_{w}\operatorname{sign}(x^{c,i})\operatorname{sign}(x^{c^{\prime},j})}{\pi\kappa(x^{c,i},x^{c^{\prime},j})}$ (112)
@@ E.2.1 Calculating $\mathbb{E}\left[\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})\right]$
$\displaystyle\begin{split}\mathbb{E}\left[\kappa(x^{c,i},x^{c,i})\right]&=\frac{1}{2}+2T(c)+\mathbb{E}[\xi_{h.o.t}]\end{split}$ (118)
@@
$\displaystyle\begin{split}S(\mu_{c},\sigma_{c})&=-\frac{1}{2}+2T(c)+\mathbb{E}[\xi_{h.o.t}],\end{split}$ (119)
which gives us:
$\displaystyle\mathbb{E}\left[\kappa(x^{c,i},x^{c,i})\right]=1+S(\mu_{c},\sigma_{c})$ (120)
@@
$\displaystyle\mathbb{E}\left[\Theta_{NTK-Erf}^{(2)}(x^{c,i},x^{c^{\prime},j})\right]$ =$\displaystyle=\begin{cases}\mathbb{E}\left[\sigma_{w}^{2}Q_{GP-Erf}^{(1)}(x^{c,i},x^{c,j})\right]+\mathbb{E}\left[\frac{4\sigma_{w}}{\pi\kappa(x^{c,i},x^{c,j})}\right]&c=c^{\prime}\\ \mathbb{E}\left[\sigma_{w}^{2}Q_{GP-Erf}^{(1)}(x^{c,i},x^{c^{\prime},j})\right]-\mathbb{E}\left[\frac{4\sigma_{w}}{\pi\kappa(x^{c,i},x^{c^{\prime},j})}\right]&c\neq c^{\prime}\end{cases},$ (143)
From (87), we know that:
To simplify the presentation, we can ignore the higher-order terms and obtain:
@@
$\displaystyle\hskip 10.0pt\approx\begin{cases}\sigma_{w}^{2}\left(1-\frac{T(c)}{2\sigma_{w}^{2}}\right)+\frac{4\sigma_{w}}{\pi}\left(\frac{3}{2}-2T(c)\right)&c=c^{\prime};i=j\\ \sigma_{w}^{2}\left(1-\frac{T(c)}{2\sigma_{w}^{2}}-\frac{T(c)^{2}}{16\sigma_{w}^{4}}\right)+\frac{4\sigma_{w}}{\pi}\left(\frac{3}{2}-\frac{T(c)^{2}}{2\sigma_{w}^{2}}-2T(c)\right),&c=c^{\prime},i\neq j\\ \sigma_{w}^{2}\left(1-\frac{T(c)+T(c^{\prime})}{4\sigma_{w}^{2}}-\frac{T(c)T(c^{\prime})}{16\sigma_{w}^{4}}\right)-\frac{4\sigma_{w}}{\pi}\left(\frac{3}{2}-\frac{T(c)T(c^{\prime})}{2\sigma_{w}^{2}}-T(c)-T(c^{\prime})\right),&c\neq c^{\prime}\\ \end{cases}$ (146)
Observe that the order of the $T(c)$ terms involved here resemble that of the NNGP scenario in (87). Thus, we can make similar conclusions regarding the role of the order of $\mu_{c},\sigma_{c}$ in determining the value of $\mathbb{E}\left[\mathcal{N}\mathcal{C}_{1}(\mathbf{H})\right]$.